OpenStudy (mbma526):

Help please will fan and medal!! 1:A rectangle has sides of 4 and 8 inches. What is the perimeter of the rectangle, expressed in feet? 2:The diameter of Mercury is 3,024 miles. Express this diameter in yards. 3:A pool can hold 1,296 gallons of water. How many quarts of water can this pool hold? 4:An elevator can hold 2,500 pounds. 5 barrels, weighing 100 pounds each, are in the elevator. About how many additional tons can the elevator hold?

OpenStudy (aaronandyson):

1.Perimeter = 2(l+b) = 2(4 + 8) = 2(12) = 24 inches 1 foot = 12 inches Therefore 24 inches = 2 feet Ans = 2 ft

OpenStudy (aaronandyson):

2) 1 mile = 1760 yards Therefore , 3025 miles = 5322240

OpenStudy (aaronandyson):

5322240 yards*

OpenStudy (mbma526):

thank you so much @AaronAndyson

OpenStudy (aaronandyson):

YW!

OpenStudy (mbma526):

:D @AaronAndyson

OpenStudy (aaronandyson):

^.^ @mbma526